# BibAgent: An Agentic Framework for Traceable Miscitation Detection in Scientific Literature

### v1

@@ A.1 Source Selection Rationale
    Article-level selection. Within the JIF-leading journal, we consider all “research article”-type items published in 2024–2025.111We exclude editorials, letters, corrigenda, and news-and-views pieces by filtering on publisher-provided article-type metadata and, when ambiguous, by manual inspection. Among these, we select the most-cited paper according to citation counts as of 2025-11-30. Citations are computed within the same JCR subject category to avoid advantaging articles that cross over into much larger neighboring fields.
@@
This procedure offers three advantages. First, it guarantees that each source paper is a field-central object around which substantial citation behavior has already accumulated, rather than a marginal or rarely cited work. Second, by choosing within each JCR category, we avoid overrepresenting large or fast-growing fields and maintain coverage across 21 high-level disciplines. Third, anchoring the time window in 2024–2025 maximizes exposure to contemporary citation practices and reduces the risk that frontier LLMs have memorized the full text, which is critical for the knowledge-blank protocol in Appendix A.2.
A.2 Knowledge-Blank Cleanroom Fact-Checking Protocol
@@
A.3 Dual-Tier Adversarial Miscitation Generation and Validation
@@
Appendix B Taxonomy Annotation Protocol
This appendix provides the full details of the expert annotation study referenced in Section 2.1. The goals of this study are twofold: (i) to test whether the five-category taxonomy is empirically usable and complete across disciplines, and (ii) to quantify how reliably independent experts can reproduce each other’s judgements when constrained by the Dependency Precedence Rule.
B.1 Sampling Strategy and Annotator Panel
Instance selection.
  1. 1.
  2. 2.
Annotator qualifications.
B.2 Annotation Materials and User Interface
  1. 1.
    presents the citing context and source excerpt side-by-side,
  2. 2.
  3. 3.
B.3 Decision Procedure and Dependency Precedence
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
B.4 Agreement Measurement
$\kappa=\frac{p_{o}-p_{e}}{1-p_{e}},$ (14)
  1. 1.
  2. 2.
B.5 Adjudication and Final Labels
The senior annotator has full access to:
  1. 1.
  2. 2.
    the three annotators’ labels, and
  3. 3.
    their free-text rationales.
@@ E.1 Acc-pass@3: Definition and Grading Pipeline
System prompt (grader).
You are an expert scientific editor.
You will be given:
1. A gold label and gold explanation for a miscitation instance.
2. A model’s predicted label and predicted explanation for the same instance.
Your task is to decide whether the model’s prediction is FULLY CORRECT.
A prediction is FULLY CORRECT only if BOTH of the following hold:
- The predicted label exactly matches the gold label.
- The predicted explanation identifies the same underlying error mechanism
  as the gold explanation (not just a generic restatement of the label).
Be strict. If the explanation is vague, only partially matches the gold
rationale, introduces incorrect reasoning, or misses key aspects of the
gold explanation, you must treat the prediction as INCORRECT.
Respond with a single token: either CORRECT or INCORRECT.
Do not output any other text.
User prompt (grader).
[INSTANCE]
Citing context:
{CITING_TEXT}
Gold label:
{GOLD_LABEL}
Gold explanation:
{GOLD_EXPLANATION}
[PREDICTION]
Predicted label:
{PRED_LABEL}
Predicted explanation:
{PRED_EXPLANATION}
Respond with exactly one token:
- CORRECT
- INCORRECT
For each instance $i$ and sample $k\in\{1,2,3\}$, the grader returns a verdict $g_{ik}\in\{\texttt{CORRECT},\texttt{INCORRECT}\}$.
@@ F.1 Protocol
We run the unmodified ICSV pipeline on the full MisciteBench-Paywall split. For every paywalled citation that admits at least one non-empty cluster, and for each backbone, we record:
@@
Figure 2: Evidence Committee behavior as a function of the number of distinct witness papers $n_{\text{voter}}$ supporting the dominant evidence statement $e^{\star}$ for a paywalled citation. Curves show non-abstention rate, conditional accuracy, and mean calibrated confidence (Appendix D); shaded bands indicate variation across backbones. The sharp and stable transition around $n_{\text{voter}}=6$ motivates the choice $K_{\min}=6$ in ICSV’s reliability-aware abstention rule.

### v2

@@ B.1 Source Selection Rationale
    Article-level selection. Within the JIF-leading journal, we consider all “research article”-type items published in 2024–2025.222We exclude editorials, letters, corrigenda, and news-and-views pieces by filtering on publisher-provided article-type metadata and, when ambiguous, by manual inspection. Among these, we select the most-cited paper according to citation counts as of 2025-11-30. Citations are computed within the same JCR subject category to avoid advantaging articles that cross over into much larger neighboring fields.
@@
This procedure offers three advantages. First, it guarantees that each source paper is a field-central object around which substantial citation behavior has already accumulated, rather than a marginal or rarely cited work. Second, by choosing within each JCR category, we avoid overrepresenting large or fast-growing fields and maintain coverage across 21 high-level disciplines. Third, anchoring the time window in 2024–2025 maximizes exposure to contemporary citation practices and reduces the risk that frontier LLMs have memorized the full text, which is critical for the knowledge-blank protocol in Appendix B.2.
B.2 Knowledge-Blank Cleanroom Fact-Checking Protocol
@@
B.3 Dual-Tier Adversarial Miscitation Generation and Validation
@@ C.4 Citation Mapping
C.5 Source Accessibility Classification (CSAC)
CSAC operates on the normalized metadata produced by the Citation Mapper and decides how each cited source will be verified. It first attempts to resolve a DOI; when a DOI is present, CSAC queries publisher and venue APIs for full text, falling back to metadata-only records when full text is restricted. When no DOI is available, CSAC constructs a metadata query from the title, author list, and year and searches publisher indices and curated open-access repositories for candidate matches. In all cases, any retrieved open-access surrogate must pass strict equivalence checks based on title similarity, author overlap, abstract content, and, when available, reference-list signatures; only then is it treated as the same scientific work as the cited source.
Based on these retrieval outcomes, CSAC assigns each reference to one of three regimes. References with verified full text—either the publisher version or an accepted open-access surrogate—are marked as accessible and routed to ACSV. References with stable metadata but no full-text access are marked as inaccessible and routed to ICSV together with their metadata snapshot (title, authors, abstract, venue, year, and any available identifiers). References for which neither DOI resolution nor repository search yields a plausible match are labeled Ghost Citations, providing a direct signal for Attribution & Traceability errors and preventing BibAgent from fabricating surrogate documents or hallucinated rationales for non-existent sources.
@@ E.1 Acc-pass@3: Definition and Grading Pipeline
System prompt (grader). User prompt (grader). For each instance $i$ and sample $k\in\{1,2,3\}$, the grader returns a verdict $g_{ik}\in\{\texttt{CORRECT},\texttt{INCORRECT}\}$.
@@ Appendix F Evidence Committee Reliability Ablation
High-signal subset used for the ablation.
For the reliability curves in Figure 3, we do not use all paywalled citations in MisciteBench-Paywall. Instead, we first identify paywalled citations $(A,B,c)$ for which the evidence statement $e^{\star}$ that is directly comparable to the citing claim $c$ admits an evidence committee of at least 25 distinct downstream open-access witness papers. Grouping by $B$ yields a pool of 62 paywalled sources. All citations included in the ablation below are drawn from this pool, and we vary the effective committee size for each such citation by subsampling witnesses within the corresponding aspect cluster.
@@ F.1 Protocol
We run the unmodified ICSV pipeline on the high-signal subset described above (62 paywalled sources and their associated citations). For every paywalled citation that admits at least one non-empty cluster, and for each backbone, we record:
@@ F.3 Choice of Threshold and Robustness
Appendix G Extended Discussion and Future Directions
While our main text focuses on the core technical contributions and empirical results, the broader implications of BibAgentand MisciteBench extend beyond the specific benchmarks we study.
First, the framework is designed to integrate into real authoring and editorial workflows rather than remain a standalone benchmark artifact. Because BibAgentoperates at the level of individual in-text citations and returns taxonomy-aligned error codes with explicit evidence bundles, it can be surfaced in draft-time assistants as a background verifier that continuously flags Citation Validity, Content Misrepresentation, Scope Extrapolation, Evidence Characterization, and Attribution & Traceability errors as authors write. The same outputs can be aggregated into paper-level integrity summaries for reviewers and editors, exposing not only how many citations are problematic but also how they fail and where manual intervention is most needed. In principle, this enables a “trust but verify” workflow in which generative drafting tools are paired with an independent auditing agent that is explicitly incentivized to disagree.
Second, the architecture of BibAgentgeneralizes naturally beyond single-language, text-only scholarship. The Document Parser & Citation Mapper can be extended to multilingual corpora by swapping in language-appropriate parsing and NLI components, while the taxonomy itself is defined at the level of evidential relations and thus does not depend on any particular disciplinary jargon. Likewise, the pipeline can be augmented to reason over non-textual artefacts—figures, tables, and code—by attaching structured evidence objects to citations (e.g., linking a claim to a specific panel in a figure or to a particular experiment in a code repository) and feeding those artefacts into the verification stages. In this view, bibliographic miscitation is one instance of a broader problem: aligning claims with heterogeneous evidence objects in a way that remains traceable and evaluable.
Third, the Evidence Committee mechanism provides a template for auditing paywalled or otherwise inaccessible material in other high-stakes settings. Grant proposals, patent applications, clinical guidelines, and policy briefs all rely on chains of references that may be partially hidden behind legal, privacy, or access constraints. In such regimes it is neither realistic nor desirable to reconstruct the protected documents themselves, but it is often possible to reconstruct what the broader community attributes to them via downstream citations, legal opinions, replication attempts, or institutional summaries. By treating these downstream artefacts as noisy committee members and insisting on reliability-aware abstention when evidence is thin or contradictory, one can extend the paywall-robust philosophy of BibAgentto domains where direct inspection is structurally impossible.
At the same time, several limitations and risks warrant emphasis. BibAgentinherits any systematic biases present in its backbones and in the citation graph itself: if certain communities or venues are systematically under-cited, their contributions will be underrepresented in the Evidence Committee, and field-normalized influence scores may inadvertently reinforce existing hierarchies. Long-document parsing and citation mapping, while robust in our experiments, can still fail in adversarially formatted or highly idiosyncratic documents, potentially leading to false negatives in miscitation detection. Moreover, making citation auditing tools widely available raises strategic concerns: sophisticated actors might attempt to game the taxonomy, crafting miscitations that sit near decision boundaries or that exploit known blind spots in NLI models and committee formation. These risks argue for conservative deployment, continuous monitoring, and independent human oversight, especially in high-stakes editorial contexts.
Despite these caveats, the central conceptual shift remains: miscitation need not be treated as an unavoidable byproduct of scale or as an anecdotal concern confined to case studies. With a principled taxonomy, a contamination-controlled benchmark, and an agentic verifier that respects both accessibility constraints and evidential uncertainty, citation integrity becomes a quantity that can be measured, improved, and enforced. Closing the loop between drafting, verification, and publication—and extending the same principles to other document genres—offers a concrete path toward research ecosystems in which claims are not only persuasive but verifiably grounded in the literature they cite.